# A Survey on Autonomous Driving Datasets: Data Statistic, Annotation, and Outlook

### v1

@@ V-A Perception Datasets
We leverage the proposed dataset evaluation metrics (see II-B) to calculate the impact scores of the collected perception datasets, subsequently selecting the top 50 datasets based on these scores to create a chronological overview, shown in Fig. 10. Concurrently, as described in III-B, we categorize datasets into onboard, V2X, drone, and others, choosing a subset from each to compile a comprehensive table of 50 datasets (Tab. II). It is worth noting that the datasets in the table are sorted by impact score within their respective categories and do not represent the overall top 50. In the following part, we choose several datasets with the highest impact scores within each sensing source and exhibit them considering their published year.
@@ V-A1 Onboard
KITTI. KITTI [13] has profoundly influenced the autonomous driving domain since its release in 2012. KITTI encompasses a variety of real-world driving scenarios recorded by various sensors, including cameras, LiDAR, and GPS/IMU. Its rich annotations and high-resolution sensor data have facilitated the development and benchmarking of numerous algorithms across a range of autonomous driving tasks, such as object detection, tracking, optical flow, depth estimation, and visual odometry.
@@
Cityscapes. Cityscapes [14] comprises a vast collection of images captured explicitly in intricate urban environments. Through meticulous labeling, Cityscapes provides pixel-level segmentation for 30 different object classes, which include various vehicle types, pedestrians, roads, and traffic sign information. Given its complexity and richness, Cityscapes has become a standard benchmark for tasks such as semantic segmentation in urban scenarios.
SYNTHIA. SYNTHIA [2] is a synthetic dataset in the realm of autonomous driving. The dataset includes 13,400 images with pixel-wise annotations for semantic segmentation. One of the standout attributes of SYNTHIA is its capability to bridge the gap between real-world and synthetic data, fostering the development of robust and transferable approaches across different domains.
Virtual KITTI. Virtual KITTI [3] crafts to closely mimic the original KITTI dataset through a virtual environment, distinguishing itself by delivering high-resolution video sequences encapsulating various traffic situations and environmental conditions. Similar to [13], Virtual KITTI supports crucial autonomous driving tasks, including object detection, semantic segmentation, and object tracking.
@@
VIPER. VIPER [5] is a synthetic dataset collected from driving, riding, and walking perspectives in a realistic virtual world, addressing the challenge of data scarcity and the high cost of labeling real-world data. VIPER contains over 250,000 video frames annotated with ground-truth data for low- and high-level vision tasks while encapsulating various weather conditions, lighting scenarios, and complex urban landscapes. Overall, VIPER presents a valuable and cost-effective tool for researchers to expedite the development of reliable and safe autonomous driving.
Apolloscapes. Apolloscapes [17] offers more than 140,000 high-resolution frames with accurate bounding boxes and pixel-wise semantic labels, critical for training and validating autonomous vehicles’ perception and navigation systems. Apolloscapes supports images and point clouds semantic segmentation, 2D/3D object detection, multi-object tracking, and lane segmentation, enabling the creation and evaluation of advanced and safe autonomous driving systems.
@@
SemanticKITTI. SemanticKITTI [6] is a notable extension of the KITTI family, focusing on semantic segmentation in the autonomous driving field. SemanticKITTI consists of over 43,000 LiDAR point cloud frames, making it one of the largest datasets for 3D semantic segmentation in outdoor environments. SemanticKITTI provides precise labels for 28 categories, such as car, road, building, etc., achieving a robust benchmark for evaluating the performance of point cloud semantic segmentation methods, underpinning numerous studies and innovations in related domains.
nuScenes. nuScenes [22] stands as an essential contribution to the field of autonomous driving, offering a rich repository of data that addresses the diverse needs of perception systems. nuScenes leverages LiDAR, radars, and cameras to record data from different urban scenes from Boston and Singapore. It is worth mentioning that its six cameras provide a comprehensive perspective of the surrounding environment, making them widely utilized in multi-view object detection tasks. In conclusion, the nuScenes dataset is a cornerstone in developing autonomous driving technologies, supporting multi-tasks and applications, and setting new benchmarks in the field.
@@
Waymo. The Waymo Open Dataset [23], introduced in 2019, significantly influences research and advancement in autonomous driving by providing an extensive size of multimodal sensory data with high-quality annotations. Key contributions of the Waymo dataset include its comprehensive coverage of driving conditions and geographics, which are pivotal for the robustness and generability of different tasks, such as detection, tracking, and segmentation.
BDD100K. BDD100K [24] dataset, released by the Berkeley DeepDrive center in 2018, is a substantial and diverse driving dataset renowned for its size and diversity. It comprises 100,000 videos, each about 40 seconds in duration. Meanwhile, it offers various annotated labels for object detection, tracking, semantic segmentation, and lane detection. This extensive compilation of data prompts advancements in the autonomous driving community, establishing itself as a challenging and versatile platform for researchers and engineers to propose and refine algorithms.
RADIATE. RADIATE [26] is the first public radar dataset, which contains 44,140 annotated frames gathered under several adversarial weather conditions, such as rain, fog, overcast, and snow. It also incorporates LiDAR and camera data, allowing full perception and understanding of the driving surroundings.
@@
Argoverse 2. As a sequel to Argoverse 1 [19], Argoverse 2 [28] introduces more diversified and complex driving scenarios, presenting the largest autonomous driving taxonomy to date. It captures various real-world driving scenarios across six cities and varying conditions. Argoverse 2 supports a wide range of essential tasks, including but not limited to 3D object detection, semantic segmentation, and tracking. In summary, the Argoverse 2 dataset offers numerous multimodal data of real-world driving scenarios, fostering innovations and advancements of algorithms and showing its substantial potential as a vital resource in autonomous driving.
@@ V-A2 V2X
V2VNet. The dataset introduced by V2VNet [78] focuses on leveraging V2V communication to allow autonomous vehicles to share information and perceive the environment from multiple viewpoints, which is crucial for detecting occluded objects and predicting the behavior of other traffic participants. This dataset was created using a high-fidelity LiDAR simulator named Lidarsim [96] that utilizes real-world data to generate realistic LiDAR point clouds for various traffic scenes. In conclusion, this work brings attention to the V2V as a promising avenue for improving the capabilities of autonomous vehicles.
DAIR-V2X. DAIR-V2X [27] is a pioneering resource in the Vehicle-Infrastructure Cooperative Autonomous Driving, providing large-scale, multi-modality, multi-view real-world data. The dataset is designed to tackle challenges such as the temporal asynchrony between vehicle and infrastructure sensors and the data transmission costs involved in such cooperative systems. The impact of the DAIR-V2X dataset on autonomous driving is significant because it sets a benchmark for complexities of vehicle-infrastructure cooperation thanks to its multiple scenarios from the real world.
Rope3D. Rope3D presented in [80] is an essential contribution to the perception systems, which addresses the critical gap in autonomous driving by leveraging data gathered from roadside cameras. Rope3D consists of 50,000 images under various environmental conditions, including different illumination (daytime, night, dusk) and weather scenarios (rainy, sunny, cloudy), ensuring a high diversity of data that reflects real-world complexities. Overall, the Rope3D dataset is a pioneering work that accelerates advancements in roadside perception for autonomous driving while being a critical tool for researchers and engineers to develop more robust and intelligent autonomous driving systems.
V2V4Real. V2V4Real [82] is the first large-scale, real-world dataset to address V2V cooperative perception. The data was collected from two vehicles with multimodal sensors like LiDAR and cameras. V2V4Real focuses on a range of perception tasks, such as cooperation 3D object detection, cooperative 3D object tracking, and Sim2Real domain adaptation. This versatility makes it an invaluable resource for developing and benchmarking autonomous driving algorithms.
@@ V-A4 Others
Pascal3D+. Pasacal3D+ [90] is an extension of the PASCAL VOC 2022 [97], overcoming the limitations of previous datasets by providing a richer and more varied set of annotations for images. Pasacal3D+ augments 12 rigid object categories, such as cars, buses, and bicycles, with 3D pose annotations and adds more images from ImageNet [98], resulting in a high degree of variability.
TT 100K. Tsinghua-Tencent 100K [15] solves the challenges of detecting and classifying traffic signs in realistic driving conditions. It provides 100K images, including 30,000 traffic-sign instances. Beyond its large data size, the high-resolution images encompass a wide range of illumination and weather conditions, making it robust for training and validating traffic sign recognition.
@@ V-B Prediction, Planning, and Control Datasets
Prediction, planning, and control datasets serve as the foundation for facilitating the training and evaluation of driving systems to forecast traffic dynamics, pedestrian movements, and other essential factors that influence driving decisions. By simulating myriad driving scenarios, they empower autonomous vehicles to make informed decisions, navigate complex environments, and maintain safety and efficiency on the road. Hence, we demonstrate in detail several high-impact datasets related to these tasks according to the data size, modalities, and citation number. We summarize the prediction, planning, and control datasets into task-specific and multi-task groups.
@@
Dr(eye)ve [53] 2018 onboard 500K frames driver’s attention prediction sunny, rainy, cloudy day, night urban, countryside, highway
@@
LOKI [109] 2021 onboard 644 scenarios TP, BP (sub-)urban
@@
TABLE III: Prediction, planning, and control datasets. We demonstrate several crucial datasets related to prediction, planning, and control. BP: behavior prediction, IP: intention prediction, MP: motion prediction, TP: trajectory prediction, MPlan: motion planning, DM: decision-making, OT: object tracking, QA: question-answering
@@ V-B1 Task-Specific Datasets
highD. The drone-based highD [18] dataset provides a large-scale collection of naturalistic vehicle trajectories on German highways, containing post-processed trajectories of 110,000 cars and trucks. The dataset aims to overcome the limitations of existing measurement methods for scenario-based safety validation, which often fail to capture the naturalistic behavior of road users or to include all relevant data with sufficient quality.
@@
PIE. The Pedestrian Intention Estimation (PIE) dataset proposed by [21] represents a significant advancement in understanding pedestrian behaviors in urban environments. It encompasses over 6 hours of driving footage recorded in downtown Toronto under various lighting conditions. The PIE dataset offers rich annotations for perception and visual reasoning, including bounding boxes with occlusion flags, crossing intention confidence, and text labels for pedestrian actions. The long continuous sequences and annotations facilitate multiple tasks like trajectory prediction and pedestrian intention prediction.
USyd. USyd [99] pushes the progress of the driver intention prediction in the context of urban intersections without traffic signals, which are common in urban settings and represent a challenge due to the lack of clear road rules and signals. The dataset incorporates over 23,000 vehicles traversing five different intersections, collected using a vehicle-mounted LiDAR-based tracking system. The data modalities included are exhaustive, providing vehicle tracks with lateral and longitudinal coordinates, heading, and velocity. This information is vital to predict driver behavior, accounting for the uncertainty inherent in human driving patterns.
@@
Argoverse. Argoverse [19] is a crucial dataset in 3D object tracking and motion forecasting. Argoverse provides ${}^{\circ}$ images from 7 cameras, forward-facing stereo imagery, and LiDAR point clouds. The recorded data covers over 300K extracted vehicle trajectories from 290km of mapped lanes. With the assistance of rich sensor data and semantic maps, Argoverse is pivotal in advancing research and development in prediction systems.
inD. The importance of the inD [100] lies in its large-scale, high-quality, and diverse set of trajectory data crucial for several applications, including road user prediction models and scenarios-based safety validation for autonomous vehicles in urban intersection environments. It covers around 11,500 different road user trajectories, e.g., vehicles, bicyclists, and pedestrians. These trajectories have a positioning error of less than 0.1 meters, which is pivotal for the reliability of the data.
PePscenes. PePscenes [101] tackles the requirement for understanding and anticipating pedestrian actions in dynamic driving surroundings. This dataset enhances the nuScenes [22] dataset by adding per-frame 2D/3D bounding box and behavior annotations, focusing on pedestrian crossing actions. One of the key attributes of [101] is incorporating various data types involving semantic maps, scene images, trajectories, and ego-vehicle states, which are necessary for creating robust models capable of understanding complex traffic scenarios.
openDD. The openDD [102] dataset focuses on analyzing and predicting traffic scenes around roundabouts, which are complex and unregulated by traffic lights. It is unique in leveraging drone-captured imagery with high resolution (4K), which spans over 62 hours of trajectory data from 501 separate flights. The dataset contains not only trajectories but also shapefiles and an extensible markup language (XML) file describing the road topology, along with a reference image for each underlying intersection.
@@
nuPlan. nuPlan [103] is the world’s first closed-loop machine learning-based planning benchmark in autonomous driving. This multimodal dataset comprises around 1,500 hours of human driving data from four cities across America and Asia, featuring different traffic patterns, such as merges, lane changes, interactions with cyclists and pedestrians, and driving in construction zones. These characters of the nuPlan dataset take into account the dynamic and interactive nature of actual driving, allowing for a more realistic evaluation.
@@
exiD. The exiD [104] trajectory dataset, presented in 2022, is a pivotal contribution to the highly interactive highway scenarios. It takes advantage of drones to record traffic without occlusion, minimizing the influence on traffic and ensuring high data quality and efficiency. This drone-based dataset surpasses previous datasets in terms of the diversity of interactions captured, especially those involving lane changes at highway entries and exits.
@@
MONA. The Munich Motion Dataset of Natural Driving (MONA) [105] is an extensive dataset, with 702K trajectories from 130 hours of videos, covering urban roads with multiple lanes, an inner-city highway stretch, and their transitions. This dataset boasts an average overall position accuracy of 0.51 meters, which exhibits the quality of the data collected using highly accurate localization and LiDAR sensors.

### v2

@@ V-A Perception Datasets
We leverage the proposed dataset evaluation metrics (see II-B) to calculate the impact scores of the collected perception datasets, subsequently selecting the top 50 datasets based on these scores to create a chronological overview, shown in Fig. 10. Concurrently, as described in III-B, we categorize datasets into onboard, V2X, drone, and others, choosing a subset from each to compile a comprehensive table of 50 datasets (Tab. II). It is worth noting that the datasets in the table are sorted by impact score within their respective categories and do not represent the overall top 50. In the following section, we choose several datasets with the highest impact scores within each sensing source and exhibit them, considering their published year.
@@ V-A1 Onboard
KITTI. KITTI [13] has been instrumental in advancing the autonomous driving domain since its release in 2012. KITTI data is recorded by various sensors, including cameras, LiDAR, and GPS/IMU, facilitating the development of algorithms for object detection, tracking, optical flow, depth estimation, and visual odometry. However, KITTI data was mainly recorded under ideal weather conditions in German urban areas. The relatively small geographic and environmental conditions scope limit its real-world applications.
@@
Cityscapes. Cityscapes [14] comprises a vast collection of images captured explicitly in intricate urban environments and has become a standard benchmark semantic segmentation. Cityscapes provides pixel-level segmentation through meticulous labeling for 30 object classes, including various vehicle types, pedestrians, roads, and traffic sign information. However, Cityscapes primarily focuses on German cities and lacks diverse weather conditions. These limitations hinder its usefulness for developing robust and generalized autonomous driving solutions.
@@
VIPER. VIPER [5] is a synthetic dataset collected based on driving, riding, and walking perspectives in a realistic virtual world. VIPER contains over 250K video frames annotated with ground-truth data for low- and high-level vision tasks. It encapsulates various weather conditions, lighting scenarios, and complex urban landscapes, making it an invaluable resource for testing the robustness of autonomous driving algorithms. Nonetheless, the transition from synthetic data to real-world application remains challenging, as algorithms must bridge the domain gap to maintain their effectiveness in real environments.
@@
SemanticKITTI. SemanticKITTI [6] consists of over 43,000 LiDAR point cloud frames, making it one of the most extensive datasets for 3D semantic segmentation in outdoor environments. SemanticKITTI provides precise labels for 28 categories, such as car, road, building, etc., achieving a robust benchmark for semantic segmentation. However, similar to the previous datasets, SemanticKITTI faces limitations in real-world applicability due to its restricted environmental diversity and geographical scope.
nuScenes. nuScenes [22] stands as an essential contribution to the field of autonomous driving, providing multimodal sensor setup, including LiDAR, radars, and cameras. It addresses the diversity in urban scenes and environmental conditions. The data recorded from Boston and Singapore cover varied driving behaviors and urban layouts, enhancing generalizability. Its six cameras provide a comprehensive perspective of the surrounding environment, making them widely utilized in multi-view object detection tasks. Yet, the dataset’s coverage of rare driving cases, such as accidents, could be expanded to better assess algorithm performance under challenging conditions.
@@
Waymo. The Waymo Open Dataset [23], introduced in 2019, significantly influences research and advancement in autonomous driving. Waymo provides an extensive size of multimodal sensory data with high-quality annotations compared to others. Key contributions of the Waymo dataset include its comprehensive coverage of driving conditions and geographics, which are pivotal for the robustness and generability of different tasks. The real-world applicability of Waymo’s data is enhanced by its diversity, although exploring its limitations in specific adverse conditions could provide deeper insights into areas needing improvement.
BDD100K. BDD100K [24] dataset is renowned for its size and diversity. It comprises 100,000 videos, each about 40 seconds in duration. Furthermore, it includes different times of day and weather conditions, offering a solid foundation for testing and improving the robustness of algorithms. Meanwhile, it provides various annotated labels for object detection, tracking, semantic segmentation, and lane detection. However, the varying quality of video annotations challenges the dataset’s real-world applicability, highlighting the need for consistent, high-quality annotations to ensure algorithm effectiveness across all conditions.
@@
Argoverse 2. As a sequel to Argoverse 1 [19], Argoverse 2 [28] introduces more diversified and complex driving scenarios, presenting the largest autonomous driving taxonomy to date. It captures various real-world driving scenarios across six cities and varying conditions. Argoverse 2 supports a wide range of essential tasks, including but not limited to 3D object detection, semantic segmentation, and tracking. The limitation of Argoverse 2 in the real world is similar to [22] and [23], considering more adversarial conditions and edge cases can extend its application field.
@@ V-A2 V2X
TUMTraf. The TUM Traffic Dataset family (TUMTraf) is a cutting-edge real-world dataset comprising 50,253 labeled frames (9,545 point clouds and 40,708 images) across five releases, capturing diverse traffic scenarios in Munich, Germany [84, 85, 86, 60]. It integrates multiple data modalities like RGB camera, event-based camera, LiDAR, GPS, and IMU. TUMTraf also provides perspectives from the infrastructure and vehicle, enabling research in cooperative perception. TUMTraf stands out for including edge cases like accidents, near-miss events, and traffic violations, providing a great resource to improve perception systems.
DAIR-V2X. DAIR-V2X [27] is a pioneering resource in the Vehicle-Infrastructure Cooperative Autonomous Driving, providing large-scale, multi-modality, multi-view real-world data. The dataset is designed to tackle challenges such as the temporal asynchrony between vehicle and infrastructure sensors and the data transmission costs involved in such cooperative systems. DAIR-V2X has been instrumental in advancing vehicle-infrastructure cooperation, setting benchmarks for addressing V2X perception tasks.
@@ V-A4 Others
Pascal3D+. Pasacal3D+ [97] is an extension of the PASCAL VOC 2022 [103], overcoming the limitations of previous datasets by providing a richer and more varied set of annotations for images. Pasacal3D+ augments 12 rigid object categories, such as cars, buses, and bicycles, with 3D pose annotations and adds more images from ImageNet [104], resulting in high variability. However, Pascal3D+ only focuses on rigid objects and may not fully address the dynamic driving environments where pedestrians and other non-rigid objects are present.
@@ V-B Prediction, Planning, and Control Datasets
Prediction, planning, and control datasets serve as the foundation for facilitating the development of driving systems. These datasets are critical for forecasting traffic dynamics, pedestrian movements, and other essential factors that influence driving decisions. Hence, we demonstrate in detail several high-impact datasets related to these tasks according to the data size, modalities, and citation number. We summarize these datasets into task-specific and multi-task groups.
TABLE III: Prediction, planning, and control datasets. We demonstrate several crucial datasets related to prediction, planning, and control. BP: behavior prediction, DAP: driver’s attention prediction, IP: intention prediction, MP: motion prediction, TP: trajectory prediction, MPlan: motion planning, DM: decision-making, OT: object tracking, QA: question-answering, DBP: driver behavior recognition
@@
Brain4Cars [105] 2015 others 2M frames Maneuver Anticipation
@@
Dr(eye)ve [59] 2018 onboard 500K frames DAP sunny, rainy, cloudy day, night urban, countryside, highway
@@
Drive&\&&Act [107] 2019 others 9.6M images DBR
DbNet [108] 2019 onboard 100 km distance DBR various road types
D2CAV [109] 2020 onboard behavioral strategy
@@
DriPE [114] 2021 others 10K pictures DBR daytime
Speak2label [115] 2021 others 586 videos DAP sunny, cloudy daytime
CoCAtt [116] 2022 onboard 11.9 hours DAP countryside
@@
Occ3D-nuScenes [119] 2024 onboard 40K frames occupancy prediction
Occ3D-Waymo [119] 2024 onboard 200K frames occupancy prediction
@@
HDD [120] 2018 onboard 104 hours data driver behavior, causal reasoning suburban, urban, highway
@@
PREVENTION [123] 2019 onboard 356 mins video IP, TP highway, urban areas
DrivingStereo [124] 2019 onboard 180K images trajectory planning sunny, rainy, cloudy daytime suburban, urban, highway
foggy, dusky elevated road, country road
@@
LOKI [126] 2021 onboard 644 scenarios TP, BP variety variety (sub-)urban
@@ V-B1 Task-Specific Datasets
highD. The drone-based highD [18] dataset provides a large-scale collection of naturalistic vehicle trajectories on German highways, containing post-processed trajectories of 110K cars and trucks. It is designed to address the shortcomings of traditional measurement techniques in scenario-based safety validation, which often miss capturing authentic road user behaviors or lack comprehensive, high-quality data. But highD is recorded under ideal weather conditions, limiting its application under adversarial weather conditions.
@@
PIE. The Pedestrian Intention Estimation (PIE) dataset proposed by [21] represents a significant advancement in understanding pedestrian behaviors in urban environments. It encompasses over 6 hours of driving footage recorded in downtown Toronto under various lighting conditions. PIE offers rich annotations for perception and visual reasoning, including bounding boxes with occlusion flags, crossing intention confidence, and text labels for pedestrian actions.
@@
Argoverse. Argoverse [19] is a crucial dataset in 3D object tracking and motion forecasting. Argoverse provides 360 images from 7 cameras, forward-facing stereo imagery, and LiDAR point clouds. The recorded data covers over 300K extracted vehicle trajectories from 290km of mapped lanes. With the assistance of rich sensor data and semantic maps, Argoverse is pivotal in advancing research and development in prediction systems. Real-world implementation has verified the effectiveness of Argoverse in urban environments, though its performance in different geographical areas may be constrained.
@@
nuPlan. nuPlan [113] is the world’s first closed-loop machine learning-based planning benchmark in autonomous driving. This multimodal dataset comprises around 1,500 hours of human driving data from four cities across America and Asia. It features different traffic patterns, such as merges, lane changes, interactions with cyclists and pedestrians, and driving in construction zones. These characters of the nuPlan dataset take into account the dynamic and interactive nature of actual driving, allowing for a more realistic evaluation. Its real-world impact is seen in developing more adaptive and context-aware planning systems.
@@
exiD. The exiD [117] trajectory dataset, presented in 2022, is a pivotal contribution to the highly interactive highway scenarios. It takes advantage of drones to record traffic without occlusion, minimizing the influence on traffic and ensuring high data quality and efficiency. This drone-based dataset surpasses previous datasets regarding the diversity of interactions captured, especially those involving lane changes at highway entries and exits. Extending the data from weather conditions and nighttime could be an improvement direction for this dataset.
@@
MONA. The Munich Motion Dataset of Natural Driving (MONA) [118] is an extensive dataset, with 702K trajectories from 130 hours of videos, covering urban roads with multiple lanes, an inner-city highway stretch, and their transitions. This dataset boasts an average overall position accuracy of 0.51 meters, which exhibits the quality of the data collected using highly accurate localization and LiDAR sensors. However, only recording the data in a particular city may limit its generalizability to other geographic locations.